OpenStudy (anonymous):

The cubic equation x^3+ax^2+bx-26=0 has 3 positive,distinct integer roots. Find the values a and b. (considering that a and b could be any value i cant understand how to find out what the integer roots are)

OpenStudy (kainui):

Hmm, well I'm trying some stuff out right now, this is an interesting question.

ganeshie8 (ganeshie8):

descartes rule of signs tells you \(a\) must be negative, and \(b\) must be positive

ganeshie8 (ganeshie8):

rational root theorem gives you the possible roots : \(\pm 1, \pm 2, \pm 13\)

Parth (parthkohli):

The product of roots is 26. By inspection, there are only three positive integral factors.

ganeshie8 (ganeshie8):

\[x^3+ax^2+bx-26 = (x-1)(x-2)(x-13)\] compare coefficients

Parth (parthkohli):

Alternative: exploit the sum of roots and cyclic sum of roots. :P

OpenStudy (epoweritheta):

x1x2x3=26 ; x1+x2+x3=-a;x1x2+x2x3+x1x3=b ; 26=1*2*13 only possible positive dinstinct integer combination , implies a=-16 ,b=41 note the problem is possible to solve only when these 3 conditions are given : positive dinstinct integer